# Can Separators Improve Chain-of-Thought Prompting?

### v2

@@ Abstract
Chain-of-thought (CoT) prompting is a simple and effective method for improving the reasoning capabilities of Large Language Models (LLMs). The basic idea of CoT is to let LLMs break down their thought processes step-by-step by putting exemplars in the input prompt. However, the densely structured prompt exemplars of CoT may cause the cognitive overload of LLMs. Inspired by human cognition, we introduce CoT-Sep, a method that strategically employs separators at the end of each exemplar in CoT prompting. These separators are designed to help the LLMs understand their thought processes better while reasoning. Interestingly, it turns out that CoT-Sep significantly improves the LLMs’ performances on complex reasoning tasks (e.g., GSM8K, AQuA, CSQA), compared with the vanilla CoT, which does not use separators. We also study the effects of the type and the location of separators tested on multiple LLMs, including GPT-3.5-Turbo, GPT-4, and LLaMA-2 7B.
1 Introduction
The use of large language models (LLMs) has significantly transformed our methods of processing information, enhancing performance across various fields. A key development in this area is Chain-of-Thought (CoT) prompting (Wei et al., 2022b), which enables LLMs to process complex reasoning. By outlining thought processes in a step-by-step manner, CoT prompting enables LLMs to demonstrate more sophisticated reasoning. By reasoning step-by-step, similar to the human cognition process, LLMs are able to tackle complicated problems with enhanced precision.
@@ 1 Introduction
In human cognition, the memory span is a fixed number of chunks, and therefore is important to organize the input sequence into chunks (Miller, 1956). Thus, chunking through strategic separations and breaks in text, such as a linebreak, plays an essential role in human comprehension and enhancing reasoning. Inspired by this psychological concept, we introduce CoT-Sep, an approach that strategically inserts separators at the optimal positions. CoT-Sep puts separators at the position where the LLM can segment information into manageable portions, therefore enabling better comprehension by the LLM. CoT-Sep leads to a improvement in the performance of LLMs over CoT prompting, underscoring the importance of structured formatting for optimizing LLM outputs. We conduct evaluations on multiple separators on arithmetic reasoning benchmarks and a commonsense reasoning benchmark, revealing that inserting separators at an adequate location and creating a pattern for LLMs to decipher outperforms the original CoT.
Various prompting methods are proposed to improve the performance of CoT and its variants (Ling et al., 2023; Long, 2023; Besta et al., 2023; Weng et al., 2023b; Zhang et al., 2023), which introduce additional modules for iterative refinement and verification of intermediate thoughts. Another line of works proposes refining input questions (Xi et al., 2023; Deng et al., 2023) to let LLMs better interpret the target reasoning tasks. However, these existing methods require multiple iterations through LLMs, leading to increased expenses and extended timeframes. Additionally, their task-specific designs make practical applications challenging. Through extensive experiments, we confirmed that CoT-Sep is straightforward and effective in improving CoT prompting, by simply introducing a separator (SEP) between exemplars.
Our empirical results on CoT-Sep indicate that using separators between exemplars in CoT prompting significantly aids the reasoning process of LLMs. For example, adding separators increases the accuracy by on AQuA dataset and by on GSM8K dataset, when tested on GPT-3.5-Turbo model. Interestingly, we discover that the proper placement of separators in prompts considerably impacts the LLM’s reasoning capabilities, potentially setting a new standard in the design of prompts for LLMs.
2 Related Work
@@
Complex tasks involving logical thinking, particularly in solving mathematical problems, are commonly challenging for natural language processing (NLP) models. (Cobbe et al., 2021; Koncel-Kedziorski et al., 2016; Patel et al., 2021). Recent progress in the development of LLMs (Touvron et al., 2023; Zhang et al., 2022; Chowdhery et al., 2022; Wei et al., 2022a; Brown et al., 2020) demonstrates remarkable capabilities in complex reasoning tasks. These works suggest that instead of directly generating final answers, LLMs perform better on reasoning tasks when guided through a step-by-step process. This approach involves using examples in few-shot settings and employing effective prompting strategies, such as the Chain-of-Thought (CoT) (Wei et al., 2023) prompting. To enhance the step-by-step reasoning process further, recent studies have employed external modules (Bostrom et al., 2022; Creswell & Shanahan, 2022; Tafjord et al., 2021; Lyu et al., 2023; Chen et al., 2023)
@@
Large Language Models (LLMs) have demonstrated impressive abilities in few-shot learning (Lu et al., 2023; Qiao et al., 2023). In particular, by incorporating a few exemplars into the prompts, LLMs can show exceptional performance without the need for finetuning on a training dataset (Weng et al., 2023a; Wang et al., 2023a), and such LLMs’ ability is termed as in-context learning (ICL). However, these approaches face challenges when encountering tasks that demand complex reasoning. This led to the emergence of CoT (Wei et al., 2022b), which is an approach that generates a series of reasoning steps to arrive at the final answer and is highly effective in solving difficult tasks. Recent work has also highlighted the in-context abilities of LLMs when combined with CoT prompting (Wang et al., 2023b; Ling et al., 2023). Building upon in-context learning, we use separators in exemplars to enhance complex reasoning.
@@
3 CoT-Sep
Consider one’s ability to process complex information. It is common to break down the information into chunks for ease of processing (Miller, 1956), and using separators can significantly enhance one’s comprehension and readability by providing visual breaks. Inspired by this observation, we introduce CoT-Sep, a method that places separators at the end of each prompt exemplar. Similar to Chain-of-Thought prompting, CoT-Sep generates a logical sequence of intermediate reasoning steps leading to the final answer. In contrast, as shown on the right of Fig. 1, when we include examples of chain-of-thought sequences into the prompt exemplars, we place separators at the end of each exemplar, which enables the LLMs to process a large amount of information in that manner.
4 Experiments
4.1 Experimental Settings
Our experiments are heavily based on the Vanilla CoT prompting (Wei et al., 2022b), and we use OpenAI’s gpt-3.5-turbo-0613111https://platform.openai.com/docs/models for the results in Table 1 and 3. For the results in Table 2, we use Meta’s llama-2-7b222https://llama.meta.com/llama2, OpenAI’s gpt-4-06131 and gpt-4-turbo, specifically the version of gpt-4-0125-preview1. For our experiments, we use benchmarks where CoT prompting leads to substantial improvement over standard prompting.
To be specific, we test on two challenging mathematical reasoning task benchmarks (GSM8K (Cobbe et al., 2021) with 1319 samples and AQuA (Ling et al., 2017) with 254 samples) and one commonsense reasoning benchmark (CSQA (Talmor et al., 2019) with 1221 samples). We follow the prompts designed by Wei et al. (2022b) and added separators to test CoT-Sep. More details regarding the exemplars used in our experiments are included in Sec. C in Appendix. We report the statistics of accuracy values obtained by running the experiments for 3 trials.
4.2 Results
@@
We also test CoT-Sep on other models including Llama2-7b, GPT-4-turbo and GPT-4, in Table 2. It turns out that compared with vanilla CoT, adding separators (either TripleSkip or TripleHash) increases the accuracy on most cases. For example, in GPT-4-turbo model, CoT-Sep (TripleSkip) enjoys 5.1% and 2.6% gap for GSM8K and AQuA datasets, respectively.
@@
4.2.1 Which Separator is Effective in Improving Reasoning?
@@
(2) TripleHash (###), (3) TripleStar (***), and two versions of HTML line break tags: (4) <br> and (5) <br/>.
Table 1 shows that although the use of various separators in CoT-Sep improves LLM reasoning compared to the vanilla CoT, TripleSkip is the most effective separator for enhancing LLM reasoning for the tested datasets, on GPT-3.5-Turbo. It is true that some separators are hurting the accuracy (e.g., inserting <br/> separator after exemplars is not desirable for GSM8K and AQuA datasets), which implies that it is necessary to appropriately choose separators.
@@
<br> and <br/>. These types of separators are placed in turn after distinct exemplars. See Table 10 for an example prompt for Heterogeneous CoT-Sep. Our result in Table 1 shows that Heterogeneous CoT-Sep outperforms vanilla CoT, implying that even in the absence of a clear choice for the best separator, CoT-Sep improves LLM reasoning compared to vanilla CoT. Thus, Heterogeneous CoT-Sep can be used in practical scenarios where we do not have any prior knowledge on the optimal separator for our new target task.
4.2.2 Does the Location/Number of Separators used in CoT-Sep Contribute to Enhancing CoT prompting?
To check the effect of the location of separators, we experiment CoT-Sep with two versions, CoT-Sep (Unit: Exemplar) and CoT-Sep (Unit: Sentence), as shown in Fig. 2. The first version adds the separator at the end of each exemplar, while the second version adds the separator between sentences within each exemplar’s CoT. Table 3 shows that CoT-Sep (Unit: Exemplar) outperforms CoT-Sep (Unit: Sentence). This can be explained by the visualization in Fig. 2: in the prompt of CoT-Sep (Unit: Sentence), the next question (say Question2) and the answer of the previous question (say Answer1) appear as a set, making it difficult for not only LLMs but also humans to understand, which results in the performance worse than CoT-Sep and even vanilla CoT. This shows that the placement of separators is crucial. Table 3 also shows the effect of the number of separators. The result indicates that our CoT-Sep (TripleSkip) method, with its specific number of three newline separators, achieves the best performance in GSM8K and AQuA benchmarks compared to the reduced number of separators. CoT-Sep (TripleSkip) is also on par with the highest performance for the CSQA benchmark.
4.2.3 Why does the Performance Gain of CoT-Sep differ by Reasoning Tasks?
As illustrated in Tables 1, 2, 3, we can observe that the performance gain of CoT-Sep, regardless of which separator is being used, varies for different reasoning tasks. We hypothesize the main reason for this variance is due to the variance of performance of CoT itself on different datasets. In particular, based on the explanation provided by (Wei et al., 2022b), we identify CoT-Sep to have the most gain when the task in question is challenging. Tables 1, 2, and 3 clearly show that the AQuA reasoning benchmark has the lowest overall accuracy, indicating that it is the most challenging. Therefore, CoT-Sep demonstrates a greater performance improvement on the AQuA benchmark compared to the other benchmarks.
| Arithmetic Reasoning | Commonsense Reasoning | ||
| Method | GSM8K | AQuA | CSQA |
| Vanilla CoT | 70.4 0.17 | 46.5 0.82 | 76.5 0.14 |
| CoT-Sep (Unit : Exemplar) | |||
[…]
| CoT-Sep (Unit : Sentence) | |||
[…]
5 Conclusion
In this paper, we introduce CoT-Sep, a method placing separators in a structured format to improve Chain-of-Thought (CoT) prompting popular in large language models. Within this framework, we place separators at the end of each prompt exemplar for better readability for LLMs. We demonstrate that CoT-Sep effectively improves CoT prompting in various LLMs across highly complex arithmetic and commonsense benchmarks through multiple experiments employing different separators and structural variations. While our experimental results are only focused on the CoT prompting scenarios, our simple idea of adding separators within the prompt can be easily plugged into various existing prompting techniques, such as Generated Knowledge Prompting (Liu et al., 2021), Self-Consistency (Wang et al., 2023b), Tree of Thought (Yao et al., 2023; Long, 2023) and GraphPrompt (Liu et al., 2023). We look forward to applying our approach to different prompting strategies in a near future.
In this section, we explain our experiment shown in Table 3 in more detail. We provide Fig. 3, which is a detailed figure of Fig. 2 with two exemplars.
Appendix B Limitations
Our research relies on CoT prompting (Wei et al., 2022b), which is based on in-context learning. Given CoT prompting, the model outputs a final answer along with a step-by-step reasoning process, allowing the generated CoT to validate the final answer. This process may cause users to excessively trust LLMs, which is critical, as LLMs may often generate incorrect outputs. This may ultimately lead to automation bias. Therefore, users should be aware of the need not to overly rely on LLMs for solving reasoning tasks or to make all decisions, despite receiving aid from them.
[…]
Due to the expense of large language models (LLMs), this study omits larger datasets and additional task categories, which may limit further insights. Furthermore, due to the limited access to computational resources, we conducted a single trial using Meta’s Llama-2-7b model, while carrying out three trials with OpenAI’s GPT models. For the same reason, we only focus on a few datasets to validate our method. We tested on two arithmetic reasoning datasets, GSM8K (Cobbe et al., 2021) and AQuA (Ling et al., 2017), as they are the most challenging arithmetic reasoning benchmarks available, and had the least impressive performance while using vanilla CoT prompting. For validation on another reasoning task, we selected the CSQA (Talmor et al., 2019) commonsense reasoning benchmark, as it is one of the most difficult commonsense reasoning benchmarks.
Appendix C Full List of Prompt Exemplars
Our prompts are primarily based on the vanilla CoT prompts provided by (Wei et al., 2022b). Full prompts for our experimental setting are included in Table 4 - 24.
| A: There are originally 3 cars. 2 more cars arrive. 3 + 2 = 5. The answer is 5. |
| A: Originally, Leah had 32 chocolates. Her sister had 42. So in total they had 32 + 42 = 74. After eating 35, they had 74 - 35 = 39. The answer is 39. |
| A: Jason started with 20 lollipops. Then he had 12 after giving some to Denny. So he gave Denny 20 - 12 = 8. The answer is 8. |
| A: Shawn started with 5 toys. If he got 2 toys each from his mom and dad, then that is 4 more toys. 5 + 4 = 9. The answer is 9. |
| A: Michael started with 58 golf balls. After losing 23 on tuesday, he had 58 - 23 = 35. After losing 2 more, he had 35 - 2 = 33 golf balls. The answer is 33. |
| A: Olivia had 23 dollars. 5 bagels for 3 dollars each will be 5 x 3 = 15 dollars. So she has 23 - 15 dollars left. 23 - 15 is 8. The answer is 8. |
[…]
| A: There are originally 3 cars. 2 more cars arrive. 3 + 2 = 5. The answer is 5. |
[…]
| A: Originally, Leah had 32 chocolates. Her sister had 42. So in total they had 32 + 42 = 74. After eating 35, they had 74 - 35 = 39. The answer is 39. |
[…]
| A: Jason started with 20 lollipops. Then he had 12 after giving some to Denny. So he gave Denny 20 - 12 = 8. The answer is 8. |
[…]
| A: Shawn started with 5 toys. If he got 2 toys each from his mom and dad, then that is 4 more toys. 5 + 4 = 9. The answer is 9. |
[…]
| A: Michael started with 58 golf balls. After losing 23 on tuesday, he had 58 - 23 = 35. After losing 2 more, he had 35 - 2 = 33 golf balls. The answer is 33. |
[…]
| A: Olivia had 23 dollars. 5 bagels for 3 dollars each will be 5 x 3 = 15 dollars. So she has 23 - 15 dollars left. 23 - 15 is 8. The answer is 8. |
[…]
| A: There are 15 trees originally. Then there were 21 trees after some more were planted. So there must have been 21 - 15 = 6. The answer is 6. *** |
| A: There are originally 3 cars. 2 more cars arrive. 3 + 2 = 5. The answer is 5. *** |
| A: Jason started with 20 lollipops. Then he had 12 after giving some to Denny. So he gave Denny 20 - 12 = 8. The answer is 8. *** |
| A: Shawn started with 5 toys. If he got 2 toys each from his mom and dad, then that is 4 more toys. 5 + 4 = 9. The answer is 9. *** |
| A: There were originally 9 computers. For each of 4 days, 5 more computers were added. So 5 * 4 = 20 computers were added. 9 + 20 is 29. The answer is 29. *** |
| A: Michael started with 58 golf balls. After losing 23 on tuesday, he had 58 - 23 = 35. After losing 2 more, he had 35 - 2 = 33 golf balls. The answer is 33. *** |
| A: There are 15 trees originally. Then there were 21 trees after some more were planted. So there must have been 21 - 15 = 6. The answer is 6. ### |
| A: Originally, Leah had 32 chocolates. Her sister had 42. So in total they had 32 + 42 = 74. After eating 35, they had 74 - 35 = 39. The answer is 39. ### |
| A: Jason started with 20 lollipops. Then he had 12 after giving some to Denny. So he gave Denny 20 - 12 = 8. The answer is 8. ### |
| A: Shawn started with 5 toys. If he got 2 toys each from his mom and dad, then that is 4 more toys. 5 + 4 = 9. The answer is 9. ### |
| A: There were originally 9 computers. For each of 4 days, 5 more computers were added. So 5 * 4 = 20 computers were added. 9 + 20 is 29. The answer is 29. ### |
| A: Olivia had 23 dollars. 5 bagels for 3 dollars each will be 5 x 3 = 15 dollars. So she has 23 - 15 dollars left. 23 - 15 is 8. The answer is 8. ### |
| A: There are 15 trees originally. Then there were 21 trees after some more were planted. So there must have been 21 - 15 = 6. The answer is 6. <br> |
| A: There are originally 3 cars. 2 more cars arrive. 3 + 2 = 5. The answer is 5. <br> |
| A: Originally, Leah had 32 chocolates. Her sister had 42. So in total they had 32 + 42 = 74. After eating 35, they had 74 - 35 = 39. The answer is 39. <br> |
| A: Shawn started with 5 toys. If he got 2 toys each from his mom and dad, then that is 4 more toys. 5 + 4 = 9. The answer is 9. <br> |
| A: There were originally 9 computers. For each of 4 days, 5 more computers were added. So 5 * 4 = 20 computers were added. 9 + 20 is 29. The answer is 29. <br> |
| A: Michael started with 58 golf balls. After losing 23 on tuesday, he had 58 - 23 = 35. After losing 2 more, he had 35 - 2 = 33 golf balls. The answer is 33. <br> |
| A: Olivia had 23 dollars. 5 bagels for 3 dollars each will be 5 x 3 = 15 dollars. So she has 23 - 15 dollars left. 23 - 15 is 8. The answer is 8. <br> |
| A: There are 15 trees originally. Then there were 21 trees after some more were planted. So there must have been 21 - 15 = 6. The answer is 6. <br/> |
| A: There are originally 3 cars. 2 more cars arrive. 3 + 2 = 5. The answer is 5. <br/> |
| A: Originally, Leah had 32 chocolates. Her sister had 42. So in total they had 32 + 42 = 74. After eating 35, they had 74 - 35 = 39. The answer is 39.<br/> |
| A: Jason started with 20 lollipops. Then he had 12 after giving some to Denny. So he gave Denny 20 - 12 = 8. The answer is 8. <br/> |
| A: There were originally 9 computers. For each of 4 days, 5 more computers were added. So 5 * 4 = 20 computers were added. 9 + 20 is 29. The answer is 29. <br/> |
| A: Michael started with 58 golf balls. After losing 23 on tuesday, he had 58 - 23 = 35. After losing 2 more, he had 35 - 2 = 33 golf balls. The answer is 33. <br/> |
| A: Olivia had 23 dollars. 5 bagels for 3 dollars each will be 5 x 3 = 15 dollars. So she has 23 - 15 dollars left. 23 - 15 is 8. The answer is 8. <br/> |
@@ Appendix C Full List of Prompt Exemplars
| is? Answer Choices: (a) 50 (b) 45 (c) 65 (d) 78 (e) 64 |

### v3

@@ Abstract
Chain-of-thought (CoT) prompting is a simple and effective method for improving the reasoning capabilities of Large Language Models (LLMs). The basic idea of CoT is to let LLMs break down their thought processes step-by-step by putting exemplars in the input prompt. However, the densely structured prompt exemplars of CoT may cause the cognitive overload of LLMs. Inspired by human cognition, we introduce CoT-Sep, a method that strategically employs separators at the end of each exemplar in CoT prompting. These separators are designed to help the LLMs understand their thought processes better while reasoning. Interestingly, it turns out that CoT-Sep significantly improves the LLMs’ performances on complex reasoning tasks (e.g., GSM8K, AQuA, CSQA), compared with the vanilla CoT, which does not use separators. We also study the effects of the type and the location of separators tested on multiple LLMs, including GPT-3.5-Turbo, GPT-4, and LLaMA-2 7B. Interestingly, the type/location of separators should be chosen appropriately to boost the reasoning capability of CoT. The code is available at https://github.com/cottonlove/CoT-SEP
Index Terms:
Chain-of-Thought (CoT), Large Language Model (LLM), LLM Reasoning, In-Context Learning¶Co-corresponding authors
I Introduction
The use of large language models (LLMs) has significantly transformed our methods of processing information, enhancing performance across various fields. A key development in this area is Chain-of-Thought (CoT) prompting [1], which enables LLMs to process complex reasoning. By outlining thought processes in a step-by-step manner, CoT prompting enables LLMs to demonstrate more sophisticated reasoning. By reasoning step-by-step, similar to the human cognition process, LLMs are able to tackle complicated problems with enhanced precision.
@@ I Introduction
In human cognition, the memory span is a fixed number of chunks, and therefore is important to organize the input sequence into chunks [2]. Thus, chunking through strategic separations and breaks in text, such as a linebreak, plays an essential role in human comprehension and enhancing reasoning. Inspired by this psychological concept, we introduce CoT-Sep, an approach that strategically inserts separators at the optimal positions. CoT-Sep puts separators at the position where the LLM can segment information into manageable portions, therefore enabling better comprehension by the LLM. CoT-Sep leads to a improvement in the performance of LLMs over CoT prompting, underscoring the importance of structured formatting for optimizing LLM outputs. We conduct evaluations on multiple separators on arithmetic reasoning benchmarks and a commonsense reasoning benchmark,
revealing that inserting separators at an adequate location and creating a pattern for LLMs to decipher outperforms the original CoT prompting technique.
Various prompting methods are proposed to improve the performance of CoT and its variants [3, 4, 5, 6, 7], which introduce additional modules for iterative refinement and verification of intermediate thoughts. Another line of works proposes refining input questions [8, 9] to let LLMs better interpret the target reasoning tasks. However, these existing methods require multiple iterations through LLMs, leading to increased expenses and extended timeframes. Additionally, their task-specific designs make practical applications challenging. Through extensive experiments, we confirmed that CoT-Sep is straightforward and effective in improving CoT prompting, by simply introducing a separator (SEP) between exemplars.
Our empirical results on CoT-Sep indicate that using separators between exemplars in CoT prompting significantly aids the reasoning process of LLMs. For example, adding separators increases the accuracy by on AQuA dataset and by on GSM8K dataset, when tested on GPT-3.5-Turbo model.
Interestingly, we discover that the proper placement of separators in prompts considerably impacts the LLM’s reasoning capabilities, potentially setting a new standard in the design of prompts for LLMs.
II Related Work
@@
Complex tasks involving logical thinking, particularly in solving mathematical problems, are commonly challenging for natural language processing (NLP) models. [10, 11, 12]. Recent progress in the development of LLMs [13, 14, 15, 16, 17] demonstrates remarkable capabilities in complex reasoning tasks. These works suggest that instead of directly generating final answers, LLMs perform better on reasoning tasks when guided through a step-by-step process. This approach involves using examples in few-shot settings and employing effective prompting strategies, such as the Chain-of-Thought (CoT) [18] prompting. To enhance the step-by-step reasoning process further, recent studies have employed external modules [19, 20, 21, 22, 23]
@@
Large Language Models (LLMs) have demonstrated impressive abilities in few-shot learning [24, 25]. In particular, by incorporating a few exemplars into the prompts, LLMs can show exceptional performance without the need for finetuning on a training dataset [26, 27], and such LLMs’ ability is termed as in-context learning (ICL). However, these approaches face challenges when encountering tasks that demand complex reasoning. This led to the emergence of CoT [1], which is an approach that generates a series of reasoning steps to arrive at the final answer and is highly effective in solving difficult tasks. Recent work has also highlighted the in-context abilities of LLMs when combined with CoT prompting [28, 3]. Building upon in-context learning, we use separators in exemplars to enhance complex reasoning.
III CoT-Sep
Consider one’s ability to process complex information. It is common to break down the information into chunks for ease of processing [2], and using separators can significantly enhance one’s comprehension and readability by providing visual breaks. Inspired by this observation, we introduce CoT-Sep, a method that places separators at the end of each prompt exemplar. Similar to Chain-of-Thought prompting, CoT-Sep generates a logical sequence of intermediate reasoning steps leading to the final answer. In contrast, as shown on the right of Fig. 1, when we include examples of chain-of-thought sequences into the prompt exemplars, we place separators at the end of each exemplar, which enables the LLMs to process a large amount of information in that manner.
IV Experiments
IV-A Experimental Settings
Our experiments are heavily based on the Vanilla CoT prompting [1], and we use OpenAI’s gpt-3.5-turbo-0613111https://platform.openai.com/docs/models for the results in Table I and IV. For the results in Table II, we use Meta’s llama-2-7b222https://llama.meta.com/llama2, OpenAI’s gpt-4-06131 and gpt-4-turbo, specifically the version of gpt-4-0125-preview1. For our experiments, we use benchmarks where CoT prompting leads to substantial improvement over standard prompting [1].
To be specific, we test on two challenging mathematical reasoning task benchmarks (GSM8K [10] with 1319 samples and AQuA [29] with 254 samples) and
one commonsense reasoning benchmark (CSQA [30] with 1221 samples). We follow the prompts designed by [1] and added separators to test CoT-Sep.
More details regarding the exemplars used in our experiments are included in https://github.com/cottonlove/CoT-SEP.
@@ IV-A Experimental Settings
We report the statistics of accuracy values obtained by running the experiments for 3 trials.
IV-B Results
@@
We also test CoT-Sep on other models including Llama2-7b, GPT-4-turbo and GPT-4, in Table II. It turns out that compared with vanilla CoT, adding separators (either TripleSkip or TripleHash) increases the accuracy on most cases. For example, in GPT-4-turbo model, CoT-Sep (TripleSkip) enjoys 5.1% and 2.6% gap for GSM8K and AQuA datasets, respectively.
@@
IV-B1 Which Separator is Effective in Improving Reasoning?
@@
Table I shows that although the use of various separators in CoT-Sep improves LLM reasoning compared to the vanilla CoT, TripleSkip is the most effective separator for enhancing LLM reasoning for the tested datasets, on GPT-3.5-Turbo. It is true that some separators are hurting the accuracy (e.g., inserting <br/> separator after exemplars is not desirable for GSM8K and AQuA datasets), which implies that it is necessary to appropriately choose separators.
(It should be noted that for the Llama2-7b model, all three trials produced identical accuracy results. Consequently, the standard deviation is not reported.)
@@
<br> and <br/>. These types of separators are placed in turn after distinct exemplars. See Table III for an example prompt for Heterogeneous CoT-Sep. Our result in Table I shows that Heterogeneous CoT-Sep outperforms vanilla CoT, implying that even in the absence of a clear choice for the best separator, CoT-Sep improves LLM reasoning compared to vanilla CoT. Thus, Heterogeneous CoT-Sep can be used in practical scenarios where we do not have any prior knowledge on the optimal separator for our new target task.
@@
IV-B2 Does the Location/Number of Separators used in CoT-Sep Contribute to Enhancing CoT prompting?
To check the effect of the location of separators,
we experiment CoT-Sep with two versions, CoT-Sep (Unit: Exemplar) and CoT-Sep (Unit: Sentence), as shown in Fig. 2. The first version adds the separator at the end of each exemplar, while the second version adds the separator between sentences within each exemplar’s CoT.
Table IV shows that CoT-Sep (Unit: Exemplar) outperforms CoT-Sep (Unit: Sentence). This can be explained by the visualization in Fig. 2:
in the prompt of CoT-Sep (Unit: Sentence),
the next question (say Question2) and the answer of the previous question (say Answer1) appear as a set, making it difficult for not only LLMs but also humans to understand, which results in the performance worse than CoT-Sep and even vanilla CoT. This shows that the placement of separators is crucial.
Table IV also shows the effect of the number of separators.
The result indicates that our CoT-Sep (TripleSkip) method, with its specific number of three newline separators, achieves the best performance in GSM8K and AQuA benchmarks compared to the reduced number of separators. CoT-Sep (TripleSkip) is also on par with the highest performance for the CSQA benchmark.
IV-B3 Why does the Performance Gain of CoT-Sep differ by Reasoning Tasks?
As illustrated in Tables I, II, IV, we can observe that the performance gain of CoT-Sep, regardless of which separator is being used, varies for different reasoning tasks. We hypothesize the main reason for this variance is due to the variance of performance of CoT itself on different datasets. In particular, based on the explanation provided by [1], we identify CoT-Sep to have the most gain when the task in question is challenging. Tables I, II, and IV clearly show that the AQuA reasoning benchmark has the lowest overall accuracy, indicating that it is the most challenging. Therefore, CoT-Sep demonstrates a greater performance improvement on the AQuA benchmark compared to the other benchmarks.
| Arithmetic Reasoning | Commonsense Reasoning | ||
| Method | GSM8K | AQuA | CSQA |
| Vanilla CoT | 70.4 0.17 | 46.5 0.82 | 76.5 0.14 |
| CoT-Sep (Unit : Exemplar) | |||
[…]
| CoT-Sep (Unit : Sentence) | |||
[…]
V Limitations
Our research relies on CoT prompting [1], which is based on in-context learning. Given CoT prompting, the model outputs a final answer along with a step-by-step reasoning process, allowing the generated CoT to validate the final answer. This process may cause users to excessively trust LLMs, which is critical, as LLMs may often generate incorrect outputs. This may ultimately lead to automation bias. Therefore, users should be aware of the need not to overly rely on LLMs for solving reasoning tasks or to make all decisions, despite receiving aid from them.
[…]
Moreover, due to the expense of large language models (LLMs), this study omits larger datasets and additional task categories, which may limit further insights. Furthermore, due to the limited access to computational resources, we conducted a single trial using Meta’s Llama-2-7b model, while carrying out three trials with OpenAI’s GPT models. For the same reason, we only focus on a few datasets to validate our method. We tested on two arithmetic reasoning datasets, GSM8K [10] and AQuA [29], as they are the most challenging arithmetic reasoning benchmarks available, and had the least impressive performance while using vanilla CoT prompting. For validation on another reasoning task, we selected the CSQA [30] commonsense reasoning benchmark, as it is one of the most difficult commonsense reasoning benchmarks.
VI Conclusion
In this paper, we introduce CoT-Sep, a method placing separators in a structured format to improve Chain-of-Thought (CoT) prompting popular in large language models. Within this framework, we place separators at the end of each prompt exemplar for better readability for LLMs. We demonstrate that CoT-Sep effectively improves CoT prompting in various LLMs across highly complex arithmetic and commonsense benchmarks through multiple experiments employing different separators and structural variations. While our experimental results are only focused on the CoT prompting scenarios, our simple idea of adding separators within the prompt can be easily plugged into various existing prompting techniques, such as Generated Knowledge Prompting [31], Self-Consistency [28], Tree of Thought [32, 4] and GraphPrompt [33]. We look forward to applying our approach to different prompting strategies and various LLMs in a near future.
In this section, we explain our experiment shown in Table IV in more detail. We provide Fig. 3, which is a detailed figure of Fig. 2 with two exemplars.
Appendix B Full List of Prompt Exemplars
Our prompts are mainly derived from the standard Chain-of-Thought (CoT) prompts provided by Wei et al. (2022). The most notable prompts used by our method, CoT-Sep, are included in the appendix. Complete prompts for our experimental setup can be found at https://github.com/cottonlove/CoT-SEP.